# online_appendix

### v2

@@ C. Concentration of the number of transactions by firm, 2023 (% of the total number of transactions)
The fourth stylized fact, illustrated by Panel A of Figure 2, is that Spain’s neighbors, Portugal and France, were the main destinations for low-value exports in 2023 (35% of the total number of export transactions). However, the ranking of export destinations is sensitive to the year analyzed. For example, the United States was the most important destination of low-value transactions in 2021 (Figure A.4 in the appendix). The difference in the ranking of countries between 2021 and 2023 will be explained when discussing the causes of the increase in the number of low-value trade transactions in the following section.
@@
China was the main source of low-value imports in 2023, accounting for 37% of the total number of import transactions. Only eight countries that appear on the list of the top 20 export destinations are also included in the list of the top 20 import origins. In particular, China, which is the most important origin of low-value imports by volume, is not included in the list of the top 20 destinations. [^12] If we added-up China and Hong-Kong, the ranking of the origins of low-value import transactions in 2021 would be similar to that in 2023 (Figure A.4 in the appendix). The ranking of export destinations is similar in 1997 and 2023 (panel B of Figure A.2 in the appendix). In contrast, China represented less than 5% of the total number of low-value import transactions in 1997. This comparison suggests that China played a key role in the growth of the number of low-value import transactions. [^13] Panel B of Figure A.3 shows that China also represented a large share of high-value import transactions in 2023. However, this share was smaller than that in low-value import transactions: 20% versus 37%.
@@
We estimated descriptive regressions to identify the country-level factors correlated with having a high share in Spain’s total number of low-value export and import transactions. Table A.1 in the appendix shows that the number of low-value export transactions is positively correlated with GDP and GDP per capita of the importer, sharing a land border and official language, and negatively correlated with distance and being a member of the EU. The number of low-value import transactions is positively correlated with the GDP of the source country, being a member of the EU, and negatively correlated with sharing an official language. Belonging to a free trade agreement does not play any role in either exports or imports. This is consistent with the fact that low-value import transactions are likely to be duty-free. [^14] Volpe-Martincus et al., (2024) get a similar result when comparing firms’ and consumers’ imports. The former would be equivalent to our high-value transactions and the latter to our low-value transactions. However, we observe that belonging to the EU has a negative effect on the share of low-value export transactions in the total number of export transactions, while it has a positive effect on imports.
@@
The fifth stylized fact, illustrated by panel B of Figure 2, is that the number of firms that only trade in low-value transactions is smaller than that that trade in high-value transactions, or in high and low-value transactions. For example, in 2023, the number of firms that only traded in low-value transactions represented 8% and 16% of all exporters and importers, respectively. These percentages were much higher than those at the beginning of the period (2% for both exporters and importers). There is an increase in the percentage of exporters that trade in large- and low-value transactions. In contrast, the percentage of this category of firms decreases among importers. We find a positive correlation between the number of large- and low-value transactions performed by a firm both for exports and imports. We also find a positive correlation between the total value of large- and low-value transactions (Table A.2 in the Appendix).
@@ 3 What explains the increase in the number of low-value trade transactions?
Since the beginning of the 1990s, some Spanish fashion firms have become internationally well known brands (Guillén,, 2005; Guillén and García-Canal,, 2010). The competitiveness of these firms has been based on the so-called fast fashion strategy, which focuses on rapidly adapting the clothing displayed in the store to changing consumer preferences (Crofton and Dopico,, 2007). In addition, these firms generate a sense of scarcity with collections that end quickly, leading consumers to purchase a garment without delay in the event that it is no longer available on the next visit to the store (Ghemawat and Nueno,, 2006). [^15] As explained by Fernandes and Tang, (2020), the fast fashion strategy may also have helped garment manufacturers in developed countries respond to low-cost competition from Asia.
@@
The fast-fashion business model requires stores to replenish frequently, a procedure that echoes the just-in-time system, introduced first by Toyota in the automotive sector and later implemented in other industries (Womack et al.,, 1990). Since all products exported by leading Spanish retailers go through a distribution center in Spain (Ghemawat and Nueno,, 2006; Rodríguez,, 2024), we expect the fast fashion strategy to generate a large number of low-value export transactions between the distribution center and stores abroad. [^16] We define as leading Spanish retailers all the brands owned by the top 2 Spanish fashion groups. We also anticipate the frequency of transactions to be higher in destinations geographically far from Spain. Because products take longer to reach their destination, firms have to respond very quickly to replenishment demands from stores to ensure that preferences do not change before goods arrive. This will require the use of air transport to serve destinations that cannot be reached by road in a few days from Spain.
@@
Using the procedure described in de Lucio et al., (2018), we identified the export transactions of the leading fashion retailers in Spain in our database. Panel A of Figure 3 shows the evolution of the number of low-value export transactions of the leading fashion retailers and all firms between 1997 and 2023. From 1997 to 2021, the increase in the number of low-value export transactions for all firms runs parallel to the increase in the number of low-value export transactions from the leading fashion retailers. This is because exports from the leading fashion retailers represented most of the volume of low-value export transactions: 63% in 2021. Panel B of Figure 3 shows that the top 10 destinations of the low-value export transactions of the leading fashion retailers in 2021 were countries that were not geographically close to Spain. [^17] The top 10 destinations accounted for 55% of the total number of low-value export transactions. In fact, countries that are closer to Spain, such as France, Germany, Italy, or Portugal, and which were some of the most important export destinations by value of the leading fashion retailers in 2021, were not included in the top 10 ranking. Furthermore, except for destinations that could be reached by road in a few days from Spain (Russian Federation, Turkey, and the UK), most of the products were transported by air. These findings are in line with the fast-fashion strategy described above. In summary, from 1997 to 2021, the leading fashion retailers were responsible for the increase in the total number of low-value export transactions. The frequency of low-value transactions, their geographical distribution, and the means used to transport the products are consistent with the fast-fashion strategy followed by these firms.
The number of low-value export transactions of the leading Spanish retailers decreases from 2021 onward. A factor contributing to this decline was the termination or reduction of operations in Russia after the invasion of Ukraine in February 2022 (de Lucio et al.,, 2024). [^18] One million export transactions to Russia from the leading fashion retailers were lost after the invasion of Ukraine. The total number of low-value export transactions decreases less than the number of low-value export transactions of the leading fashion retailers. This indicates that firms not related to the fashion industry began to account for a larger share of low-value export transactions. To explain this fact, we need to introduce the second explanation for the increase in the number of low-value transactions. We do it analyzing, first, the evolution of the number of low-value import transactions.
@@
A. Number of low-value export transactions from the leading fashion retailers and from all firms, 1997-2023
B. Top 10 destinations by volume for low-value export transactions from leading fashion retailers, 2021
@@ C. Number and share of low-value transactions in Spain’s imports from China, 1997-2023
This staggering increase coincides with the appearance of online retail platforms founded in China, such as Shein or Temu, which have become very attractive to consumers due to their high promotions and discounts. [^19] For example, in August 2023, Temu and Shein were the first and third most downloaded applications in Spain (see “Temu, Shein, Miravia: el agresivo comercio electrónico chino que rivaliza con Amazon” by Pablo G. Bejarano, www.elpais.com, 11 November 2023 at https://elpais.com/tecnologia/2023-11-11/temu-shein-miravia-el-agresivo-comercio-electronico-chino-que-rivaliza-con-amazon.html. These online platforms allow consumers to access a wide range of low-value Chinese products while providing Chinese manufacturers with access to a global market. Transactions mediated through these platforms are denoted as direct to consumer, because purchases go directly from the overseas producer to the local consumer (Fajgelbaum and Khandelwal,, 2024).
@@
Our data set does not provide information to determine whether a transaction is the result of an online or an offline purchase. Having this limitation in mind, we examine whether other facts are consistent with the direct-to-consumer explanation for the rise in the number of low-value import transactions from China from 2020 onward. In 2023, a firm concentrated almost half of low-value import transactions originating in China. This firm did not receive imports from China in a specific province in Spain, but they were scattered throughout all Spanish provinces. Furthermore, this firm recorded 300,000 different import transactions from China in each monthly customs declaration. This very large number of declared transactions is consistent with the fact that online platforms, or their logistic partners, can fulfill the customs declaration on behalf of their final customers, when the transaction is equal or below the EU de minimis threshold. It is very unlikely that a firm would fraction imports for its own use into so many low-value transactions to avoid the payment of tariffs, because such a large number of transactions would raise the suspicion of tax authorities. Therefore, the geographical distribution and the large number of transactions are consistent with the argument that products were sent directly from Chinese producers to Spanish customers and that the importer was an operator that facilitated customs clearance and shipments.
The top importer by number of low-value transactions from China in 2023 was the same as in 2022. However, before 2022, the main importers were a large retailer and two fashion firms. They firms received all low-value imports from China in the province where their headquarters and main warehouse were located. We do not have information to determine whether the increase in the number of low-value import transactions from China of these firms was explained by the expansion of their e-commerce sales or the use of a just-in-time strategy. However, there are two facts that are consistent with the argument that these purchases were likely to be linked to these firms’ e-commerce activities.
@@
First, in the case of the large retailer, the number of low-value import transactions from China was almost negligible in 2019. However, it became the importer with the largest number of low-value transactions from China in 2020. That year, due to Covid, non-essential stores remained closed for some weeks. Furthermore, even when the stores opened after the strictest confinement measures were lifted, the shopping experience became less pleasant for the customers due to health-related restrictions. In parallel, the retailer reported a very large increase in e-commerce sales between 2019 and 2020. These facts suggest that it was the increase in purchases on this firm’s online marketplace that contributed to the increase in the number of low-value import transactions from China. Second, as argued above, the fact that the large retailer and the two fashion firms reported very large numbers of low-value import transactions in their customs declarations is consistent with the documentation of e-commerce-based transactions.
@@
We find that the four firms that explain the rise in the number of low-value import transactions from China are also responsible for one third of the total increase in the number of low-value import transactions between 1997 and 2023. The remaining 100,000 firms that performed low-value import transactions each contributed a very tiny fraction to the remaining two-thirds of the increase. [^20] The exceptions are two firms that represented 4% and 2% of the increase that we could not identify by name. Both firms specialize in low-value imports from Germany.
@@ 4 Conclusion
The increase in the share of low-value transactions can have important implications on international trade. Online platforms have allowed consumers, particularly those with low income, to have a more direct involvement in trade, providing them with a greater leverage in shaping trade flows. In addition, the increase in low-value import transactions may lead countries to use the de minimis threshold as a trade-restriction measure.
Appendix for “The increase in the number of low-value transactions in international trade”
This appendix displays additional Tables and Figures mentioned in the main text.
Additional tables
Table A.1: Estimates from descriptive regressions analyzing the factors affecting a partner’s share in the total number of low-value transactions
(1) (2)
Exports Imports
GDP (ln) 0.548a 0.858a
(0.057) (0.283)
GDP pc (ln) 0.400a -0.113
(0.093) (0.451)
Distance (ln) -0.655a 0.370
(0.125) (0.426)
Contiguity 1.362b 0.806
(0.565) (0.615)
Language 1.007a -2.064a
(0.284) (0.591)
European Union -1.474a 2.276a
(0.274) (0.595)
Free trade agreement 0.019 -0.607
(0.239) (0.912)
Observations 192 185
Pseudo-R2 0.668 0.597
Note: Data is for 2021. In columns 1 and 2 the dependent variable is the share of a country in Spanish total number of low-value export and import transactions, respectively. We estimate a Poisson pseudo-maximum likelihood model. Distance is the harmonic mean of the population-weighted distance between the most populated cities. Contiguity is an indicator variable that turns one if the partner country shares a land border with Spain. Language is an indicator variable that turns one if Spain and the partner country share the same official language. European Union is an indicator variable that turns one if the partner country belongs to the European Union. Free trade agreement is an indicator variable that turns one if Spain and the partner country belong to the same free trade agreement. GDP and GDP per capita data are from World Bank’s World Development Indicators database. The remaining independent variables are from the CEPII gravity database. Standard errors are in parentheses. a, b, and c: statistically significant at 1%, 5%, and 10%, respectively.
Table A.2: Correlation between large- and low-value transactions at the firm level
Number Value
(1) (2) (3) (4)
Exports Imports Exports Imports
Large valueft 0.43080a 19.69778a 0.00016a 0.00060a
(0.06503) (1.84782) (0.00005) (0.00019)
Observations 1504377 2760297 1504377 2760297
Pseudo-R2 0.72290 0.54298 0.83725 0.67368
[…]
Additional figures
Figure A.1: Robustness analyses on the share of low-value transactions in the total number of trade transactions

### v3

@@ C. Concentration of the number of transactions by firm, 2023 (% of the total number of transactions)
The fourth stylized fact, illustrated by Panel A of Figure 2, is that Spain’s neighbors, Portugal and France, were the main destinations for low-value exports in 2023 (35% of the total number of export transactions). However, the ranking of export destinations is sensitive to the year analyzed. For example, the United States was the most important destination of low-value transactions in 2021 (Figure A.4 in the Online Appendix). The difference in the ranking of countries between 2021 and 2023 will be explained when discussing the causes of the increase in the number of low-value trade transactions in the following section.
@@
China was the main source of low-value import transactions in 2023, accounting for 37% of the total number of import transactions. Only eight countries that appear on the list of the top 20 export destinations are also included in the list of the top 20 import origins. In particular, China, which is the most important origin of low-value imports by volume, is not included in the list of the top 20 destinations. [^10] If we added-up China and Hong-Kong, the ranking of the origins of low-value import transactions in 2021 would be similar to that in 2023 (Figure A.4 in the Online Appendix). The ranking of export destinations is similar in 1997 and 2023 (panel B of Figure A.2 in the Online Appendix). In contrast, China represented less than 5% of the total number of low-value import transactions in 1997. This comparison suggests that China played a key role in the growth of the number of low-value import transactions. [^11] Panel B of Figure A.3 shows that China also represented a large share of high-value import transactions in 2023. However, this share was smaller than that in low-value import transactions: 20% versus 37%.
@@
We estimated descriptive regressions to identify the country-level factors correlated with having a high share in Spain’s total number of low-value export and import transactions. Table A.1 in the Online Appendix shows that the number of low-value export transactions is positively correlated with GDP and GDP per capita of the importer, sharing a land border and official language, and negatively correlated with distance and being a member of the EU. The number of low-value import transactions is positively correlated with the GDP of the source country, being a member of the EU, and negatively correlated with sharing an official language. Belonging to a free trade agreement does not play any role in either exports or imports. This is consistent with the fact that low-value import transactions are likely to be duty-free. [^12] Volpe-Martincus et al., (2024) get a similar result when comparing firms’ and consumers’ imports. The former would be equivalent to our high-value transactions and the latter to our low-value transactions. However, we observe that belonging to the EU has a negative effect on the share of low-value export transactions in the total number of export transactions, while it has a positive effect on imports.
@@
The fifth stylized fact, illustrated by panel B of Figure 2, is that the number of firms that only trade in low-value transactions is smaller than that that trade in high-value transactions, or in high and low-value transactions. For example, in 2023, the number of firms that only traded in low-value transactions represented 8% and 16% of all exporters and importers, respectively. These percentages were much higher than those at the beginning of the period (2% for both exporters and importers). There is an increase in the percentage of exporters that trade in large- and low-value transactions. In contrast, the percentage of this category of firms decreases among importers. We find a positive correlation between the number of large- and low-value transactions performed by a firm both for exports and imports. We also find a positive correlation between the total value of large- and low-value transactions (Table A.2 in the Online Appendix).
@@ 2 What explains the increase in the number of low-value trade transactions?
Since the beginning of the 1990s, some Spanish fashion firms have become internationally well known brands (Guillén,, 2005; Guillén and García-Canal,, 2010). The competitiveness of these firms has been based on the so-called fast fashion strategy, which focuses on rapidly adapting the clothing displayed in the store to changing consumer preferences (Crofton and Dopico,, 2007). In addition, these firms generate a sense of scarcity with collections that end quickly, creating a fear of missing out feeling among customers, which leads them to purchase a garment without delay in the event that it is no longer available on the next visit to the store (Ghemawat and Nueno,, 2006). [^13] As explained by Fernandes and Tang, (2020), the fast fashion strategy may also have helped garment manufacturers in developed countries respond to low-cost competition from Asia.
@@
The fast-fashion business model requires stores to replenish frequently, a procedure that echoes the just-in-time system, introduced first by Toyota in the automotive sector and later implemented in other industries (Womack et al.,, 1990). Since all products exported by leading Spanish retailers go through a distribution center in Spain (Ghemawat and Nueno,, 2006; Rodríguez,, 2024), we expect the fast-fashion strategy to generate a large number of low-value export transactions between the distribution center and stores abroad. We also anticipate the frequency of transactions to be higher in destinations geographically far from Spain. Because products take longer to reach their destination, firms have to respond very quickly to replenishment demands from stores to ensure that preferences do not change before goods arrive. This will require the use of air transport to serve destinations that cannot be reached by road in a few days from Spain.
@@
Using the procedure described in de Lucio et al., (2018), we identified the export transactions of the leading fashion retailers in Spain in our database. Panel A of Figure 3 shows the evolution of the number of low-value export transactions of the leading fashion retailers and all firms between 1997 and 2023. From 1997 to 2021, the increase in the number of low-value export transactions for all firms runs parallel to the increase in the number of low-value export transactions from the leading fashion retailers. This is because exports from the leading fashion retailers represented most of the volume of low-value export transactions: 63% in 2021. Panel B of Figure 3 shows that the top 10 destinations by volume of low-value export transactions of the leading fashion retailers in 2021 were countries that were not geographically close to Spain. [^14] The top 10 destinations accounted for 55% of the total number of low-value export transactions. In fact, countries that are closer to Spain, such as France, Germany, Italy, or Portugal, and which were some of the most important export destinations by value of the leading fashion retailers in 2021, were not included in the top 10 ranking of low-value export transaction destinations. Furthermore, except for destinations that could be reached by road in a few days from Spain (Russian Federation, Turkey, and the UK), most of the products were transported by air. These findings are in line with the fast-fashion strategy described above. In summary, from 1997 to 2021, the leading fashion retailers were responsible for the increase in the total number of low-value export transactions. The frequency of low-value transactions, their geographical distribution, and the means used to transport the products are consistent with the fast-fashion strategy followed by these firms.
@@
A. Number of low-value export transactions from the leading fashion retailers and from other firms, 1997-2023
B. Top 10 destinations by volume of low-value export transactions from leading fashion retailers, 2021 (millions)
@@ C. Number and share of low-value transactions in Spain’s imports from China, 1997-2023
The number of low-value export transactions of the leading Spanish retailers decreases from 2021 onward. A factor contributing to this decline was the termination or reduction of operations in Russia after the invasion of Ukraine in February 2022 (de Lucio et al.,, 2024). [^15] Export transactions from the leading fashion retailers to Russia decreased in one million after the invasion of Ukraine. Furthermore, the total number of low-value export transactions decreases less than the number of low-value export transactions of the leading fashion retailers. This indicates that firms not related to the fashion industry began to account for a larger share of low-value export transactions. To explain this fact, we need to introduce the second explanation for the increase in the number of low-value transactions. We do it analyzing the evolution of low-value transactions in imports.
@@
This staggering increase coincides with the popularization of online retail platforms founded in China, such as Shein or Temu. [^16] For example, in August 2023, Temu and Shein were the first and third most downloaded applications in Spain (see “Temu, Shein, Miravia: el agresivo comercio electrónico chino que rivaliza con Amazon” by Pablo G. Bejarano, www.elpais.com, 11 November 2023 at https://elpais.com/tecnologia/2023-11-11/temu-shein-miravia-el-agresivo-comercio-electronico-chino-que-rivaliza-con-amazon.html. These online platforms allow consumers to access a wide range of low-value Chinese products while providing Chinese manufacturers with access to a global market. They follow a strategy that resembles an extreme version of the fast-fashion strategy, introducing thousands of new products every day and offering “cannot-miss” promotions and discounts. Transactions mediated through these platforms are denoted as direct to the consumer because purchases go directly from the overseas producer to the local consumer.
@@
Our data set does not provide information on whether a transaction is the result of an online or an offline purchase. Having this limitation in mind, we examine whether other facts are consistent with the direct-to-consumer explanation for the rise in the number of low-value import transactions from China from 2020 onward. In 2023, a firm concentrated almost half of low-value import transactions originating in China. This firm did not receive imports from China in a specific province in Spain, but they were scattered throughout all Spanish provinces. Furthermore, this firm recorded 300,000 different import transactions from China in each monthly customs declaration. This very large number of declared transactions is consistent with the fact that online platforms, or their logistic partners, can fulfill the customs declaration on behalf of their final customers, when the transaction is equal or below the EU de minimis threshold. It is very unlikely that a firm would fraction imports for its own use into so many low-value transactions to avoid the payment of tariffs, because such a large number of transactions would raise the suspicion of tax authorities. Therefore, the geographical distribution and the large number of transactions are consistent with the argument that products were sent directly from Chinese producers to Spanish customers and that the importer was an operator that facilitated customs clearance and shipments.
@@
The top importer by number of low-value transactions from China in 2023 was the same as in 2022. However, before 2022, the main importers were a large retailer and two fashion firms. These firms received all low-value imports from China in the province where their headquarters and main warehouse were located. We do not have information to determine whether the increase in the number of low-value import transactions from China of these firms was explained by the expansion of their E-commerce sales or the use of a just-in-time strategy. However, there are two facts that are consistent with the argument that these purchases were likely to be linked to these firms’ E-commerce activities. First, in the case of the large retailer, the number of low-value import transactions from China was almost negligible in 2019. However, it became the importer with the largest number of low-value transactions from China in 2020. That year, due to Covid, non-essential stores remained closed for some weeks. Furthermore, even when the stores opened after the strictest confinement measures were lifted, the shopping experience became less pleasant for the customers due to health-related restrictions. In parallel, the retailer reported a very large increase in E-commerce sales between 2019 and 2020. These facts suggest that it was the increase in purchases on this firm’s online marketplace that contributed to the increase in the number of low-value import transactions from China. Second, as argued above, the fact that the large retailer and the two fashion firms reported very large numbers of low-value import transactions in their customs declarations is consistent with the documentation of low-value E-commerce-based transactions.
@@
We find that the four firms that explain the rise in the number of low-value import transactions from China are also responsible for one third of the total increase in the number of low-value import transactions between 1997 and 2023. The remaining 100,000 firms that performed low-value import transactions each contributed a very tiny fraction to the remaining two-thirds of the increase. [^17] The exceptions are two firms that represented 4% and 2% of the increase and specialize in low-value imports from Germany.
@@ 3 Conclusion
The increase in the share of low-value transactions can have important implications on international trade. Online platforms have allowed consumers, particularly those with low income, to have a more direct involvement in trade, providing them with a greater leverage in shaping trade flows. In addition, the increase in low-value import transactions may lead countries to adopt protection measures by reducing the de minimis threshold. Finally, it has also raised concerns about the environmental impact of trade and data protection.